OpenStudy (anonymous):

Can someone teach me how to do this? Thanks in advance! and Please hurry! Simplify 2s^2 - 5s - 12 / 2s^2 - 9s + 4

OpenStudy (jhannybean):

take one part at a time, let's start with 2s^2 - 5s - 12 multiply the leading coefficient to the end term. s^2 - 5s - 24 what two numbers multiply to give -24, but add to give -5?

OpenStudy (anonymous):

-8 and 3

OpenStudy (jhannybean):

good job :)

OpenStudy (jhannybean):

Now we're going to factor s^2 -5s - 24 using -8 and 3 (s - 8)(s + 3) with me so far?

OpenStudy (anonymous):

Alright

OpenStudy (jhannybean):

now we're going to divide these factors by the leading coefficient we used to multiply them in order to get the factors ( s - 8/2)(s+3/2) we're going to reduce these two factors as much as possible, and if we cant reduce them, we're going to take the denominator and multiply it to the variable. Like so (s - 4)(2s + 3) See what i did there?

OpenStudy (anonymous):

Hmm, Where did the 2 in (2s + 3) come from?

OpenStudy (jhannybean):

we had (s + 3/2) and because we couldnt reduce it, the other simple form we could put it in was (2s + 3) by multiplying "2" to the "s"

OpenStudy (anonymous):

Ok I understand

OpenStudy (jhannybean):

Awesome. Now the first part is simplified.

OpenStudy (anonymous):

ok

OpenStudy (jhannybean):

Now can you work 2s^2 - 9s + 4 the same way we did the first one? I will guide you through it :)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So s^2 - 9s + 8 then -1 and -8 so (s - 1) and (s - 8) am I doing ok so far?

OpenStudy (jhannybean):

you're doing absolutely fine :D Now we're going to divide by the leading coefficient just like the last one. what would you get?

OpenStudy (anonymous):

(2s - 1) and (s - 4)

OpenStudy (jhannybean):

yay!! you've got this!

OpenStudy (anonymous):

haha

OpenStudy (jhannybean):

now we're going to put it all together. 2s^2 -5s - 12 = (s - 4)(2s + 3) 2s^2 - 9s + 4 = (s - 4)(2s - 1) right?

OpenStudy (anonymous):

ok, now we can get rid of both (s - 4) s right? making (2s + 3) over (2s - 1) right?

OpenStudy (jhannybean):

Yep :D You got it.

OpenStudy (jhannybean):

And you've got your answer ^_^

OpenStudy (anonymous):

Yay! Thanks very much! Got time for one more problem?

OpenStudy (jhannybean):

I don't actually x_x heading down to dinner, haha xD

OpenStudy (anonymous):

Ok thanks again have fun! Enjoy your dinner!

OpenStudy (jhannybean):

Thanks :D